OpenStudy (anonymous):

square root of 24x^8y^4? show work? thank you

myininaya (myininaya):

\[\sqrt{24x^8y^4}=\sqrt{24}\sqrt{x^8}\sqrt{y^4}=\sqrt{24}x^4y^2=\sqrt{4\cdot 6}x^4y^2=2\sqrt{6}x^4y^2\] \[=2x^4y^2\sqrt{6}\]

OpenStudy (anonymous):

Wow thank you, You're amazing! =D

myininaya (myininaya):

thanks :)